Examine which of the following conicoids are central and which are non-central. Also determine which of the central conicoids have centre at the origin.

(i) π‘₯2 + 𝑦2 + 𝑧2 + 4π‘₯ + 3𝑦 βˆ’ 𝑧 = 0

(ii) 2π‘₯2 βˆ’ 𝑦2 βˆ’ 𝑧2 + π‘₯𝑦 + 𝑦𝑧 βˆ’ 𝑧π‘₯ = 1

(iii) π‘₯2 + 𝑦2 βˆ’ 𝑧2 βˆ’ 2π‘₯𝑦 βˆ’ 3𝑦𝑧 βˆ’ 6𝑧π‘₯ + π‘₯ βˆ’ 2𝑦 + 5𝑧 + 4 = 0


1
Expert's answer
2022-02-14T12:04:19-0500

A conicoid given by equation

"ax^2 +by^2 +cz^2 +2fyz+2gxz+2hxy"




"+2ux+2vy+2wz+d=0"

has a point "(x_0 ,y_0 ,z_0 )" as a center if


"\begin{cases}
ax_0+hy_0+gz_0+u=0 \\
hx_0+by_0+fz_0+v=0 \\
gx_0+fy_0+cz_0+w=0
\end{cases}"



(i)

"a=1, b=1, c=1,"

"f=g=h=0,"

"u=2,v=3\/2, w=-1\/2"

Then


"\begin{cases}
x_0+2=0 \\
y_0+3/2=0 \\
z_0-1/2=0
\end{cases}"

The central conicoid has a center "(-2,-3\/2,1\/2)."


(ii)

"a=2, b=-1, c=-1,"

"f=1\/2,g=-1\/2,h=1\/2,"

"u=v=w=0"

Then


"\begin{cases}
2x_0+\dfrac{1}{2}y_0-\dfrac{1}{2}z_0=0 \\ \\
\dfrac{1}{2}x_0-y_0+\dfrac{1}{2}z_0=0 \\ \\
-\dfrac{1}{2}x_0+\dfrac{1}{2}y_0-z_0=0
\end{cases}"

The central conicoid has a center "(0,0,0)."


(iii)

"a=1, b=1, c=-1,"

"f=-3\/2,g=-3,h=-1,"

"u=1\/2,v=-1,w=5\/2"

Then


"\begin{cases}
x_0-y_0-3z_0+\dfrac{1}{2}=0 \\ \\
-x_0+y_0-\dfrac{3}{2}z_0-1=0 \\ \\
-3x_0-\dfrac{3}{2}y_0-z_0+\dfrac{5}{2}=0
\end{cases}"

The central conicoid has a center "(\dfrac{49}{162},\dfrac{92}{81},-\dfrac{1}{9})."